Find by lagrange’s formula, the value of
U4 if U3 = 16 , U5 = 36 , U7 = 64 , U8 = 81 and U9 = 100
1
Expert's answer
2020-12-28T18:47:31-0500

Here the intervals are unequal.


"\begin{matrix}
 x_0=3, & x_1=5, & x_2=7, & x_3=8, & x_4=9 \\
 y_0=16, & y_1=36, & y_2=64, & y_3=81,& y_4=100
\end{matrix}"

"y=f(x)=(x+1)^2"

Put "x=4"


"f(4)=(4+1)^2=25"

By Lagrange’s interpolation formula we have

"y=f(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)(x-x_4)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)(x_0-x_4)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)(x-x_4)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)(x_1-x_4)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)(x-x_4)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)(x_2-x_4)}\times y_2"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)(x-x_4)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)(x_3-x_4)}\times y_3"


"+\dfrac{(x-x_0)(x-x_1)(x-x_2)(x-x_3)}{(x_4-x_0)(x_4-x_1)(x_4-x_2)(x_4-x_3)}\times y_4"



Put "x=4"


"f(4)=\dfrac{(4-5)(4-7)(4-8)(4-9)}{(3-5)(3-7)(3-8)(3-9)}\times 16"

"+\dfrac{(4-3)(4-7)(4-8)(4-9)}{(5-3)(5-7)(5-8)(5-9)}\times 36"

"+\dfrac{(4-3)(4-5)(4-8)(4-9)}{(7-3)(7-5)(7-8)(7-9)}\times 64"

"+\dfrac{(4-3)(4-5)(4-7)(4-9)}{(8-3)(8-5)(8-7)(8-9)}\times 81"

"+\dfrac{(4-3)(4-5)(4-7)(4-8)}{(9-3)(9-5)(9-7)(9-8)}\times 100"

"=25"

"U4=25"